Eleven, Lucas, Chief Hopper, The Demogorgon, and Karen have all been invited to Ted Wheeler’s strange dinner party. They arrive randomly, and each individual arrives at a different time. What is the probability that Chief Hopper will arrive first, Karen will be the second to arrive, and Eleven will arrive last?
Answer #1

Eleven, Lucas, Chief Hopper, The Demogorgon, and Karen 5 people are being invited

Total number of ways these 5 people can arrive = 5! = 5*4*3*2*1 = 120

Number of ways Chief Hopper will arrive first, Karen will be the second to arrive, and Eleven will arrive last = 2 ( fixed places for 3 and rest 2 can take anyplace either 3rd or 4th )

P[ Chief Hopper will arrive first, Karen will be the second to arrive, and Eleven will arrive last ] = Number of ways Chief Hopper will arrive first, Karen will be the second to arrive, and Eleven will arrive last / Total number of ways these 5 people can arrive

P[ Total number of ways these 5 people can arrive ] = 2/120

P[ Total number of ways these 5 people can arrive ] = 1/60